OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6. Answer y-1=6(x+5) y+5=6(x-1) y+1=6(x-5) y-5=6(x+1)

OpenStudy (lgbasallote):

do you have an answer?

OpenStudy (anonymous):

? No?

OpenStudy (anonymous):

the answers are up there i dont know which one is right though haha

OpenStudy (lgbasallote):

point slope form is \[\Huge y - y_1 = m (x-x_1)\]

OpenStudy (lgbasallote):

x1 is the x coordinate y1 is the y coordinate m is the slope

OpenStudy (lgbasallote):

do you still need more help?

OpenStudy (anonymous):

yeah ha

OpenStudy (lgbasallote):

try substituting..what's the x coordinate in the problem?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

??

OpenStudy (lgbasallote):

correct. what's the y-coordinate?

OpenStudy (anonymous):

-1

OpenStudy (lgbasallote):

and the slope?

OpenStudy (anonymous):

6

OpenStudy (lgbasallote):

good. so now what is the point slope form?

OpenStudy (anonymous):

would M be the slope?

OpenStudy (anonymous):

would it be y+5=6(x-1)

OpenStudy (lgbasallote):

nope...remember y - y1 = 6(x-x1) what you did is y - x1 = 6(x-y1)

OpenStudy (lgbasallote):

5 is x coordinate -1 is y coordinate

OpenStudy (anonymous):

OH! so its y-1=6(x+5)